OpenStudy (firejay5):

I just took a quiz over Dividing Rational Numbers and this question I believe is wrong or something, and I need someone's opinion on this. Divide. 3.4 ÷ 0.05 A. 65 B. 54 C. 6.3 D. 63 *It says the answer is D, but I got 68 as my answer, so either I am wrong or the question on the quiz is wrong. Help me and explain process for my understanding.

OpenStudy (jiteshmeghwal9):

U r right

OpenStudy (firejay5):

I just think it's a glitch or something lol

OpenStudy (daniel_chernioglo):

Yup, i got the 68 as the answer as well

OpenStudy (jiteshmeghwal9):

3.4/0.05 =(34/10)×(100/5) =(34×10)/5 =68

OpenStudy (firejay5):

It's from SoftSchools.com

OpenStudy (firejay5):

It's from SoftSchools.com